Circular No. 7892

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS M. Orio, National Institute for Astrophysics, Torino, and University of

Wisconsin; D. Harbeck, Max-Planck-Institut für Astronomie, Heidelberg; and J. Gallagher and C. Woodward, University of Wisconsin, report that images (0′′.141/pixel) of the “light echo” around V838 Mon (IAUC 7859), taken with the WIYN 3.5-m telescope on Apr. 19 (R band, and Hα narrow filter) and 21 (V band, and [O III] narrow filter), show a filled, approximately circular shell, not limb brightened, with a diameter of 35′′ on both nights. They add: “The contour of the shell is irregular, with a broad structure in the shape of an arc and clumps but no small-scale structure. There is more structure in V than in R. Lack of conspicuous Hα and [O III] line emission confirms that it is a reflection, rather than an ionization, nebula. An R image obtained on Apr. 30 at the WIYN 0.90-m telescope (scale 0′′.60/pixel) still shows a 35′′ diameter, so the outer edge of the reflecting material was reached before Apr. 22.”

H. E. Bond, N. Panagia, and W. B. Sparks, Space Telescope Science Institute; S. G. Starrfield, Arizona State University; and R. M. Wagner, University of Arizona, report that the rapidly evolving light echo around V838 Mon was imaged with the Hubble Space Telescope on Apr. 30. The newly installed Advanced Camera for Surveys was used in its polarimetric mode with a B filter. The echo shows significant linear polarization, as expected for scattering off interstellar grains, but numerical results await forthcoming calibration observations for the ACS. The structure of the echo is complex: V838 Mon is surrounded by at least 11 separate, approximately circular concentric rings or arcs, out to a radius of 22′′. Most of these arcs correspond to the intersections of the light-echo paraboloid with sheets of interstellar material in the line-of-sight (similar to the transient echo rings seen around SN 1987A), but it is possible that some of the features at the smallest radii arise in circumstellar material from hypothetical previous outbursts of V838 Mon. The echo also contains several highly structured, non-concentric filamentary features. The prominence of the echo strongly suggests that V838 Mon was intrinsically extremely luminous during its prolonged outburst, and is thus unrelated to the fainter FG Sge-like ‘bornagain’ red giants. It may instead be related to the very luminous extragalactic red variable ‘M31 RV’ that had a long outburst in 1988. Analysis of the polarimetric images will yield geometric constraints on the distance of V838 Mon. An additional HST observation will be made during the week starting May 20. Continued intensive ground-based monitoring of the light curve and the light echo is urged.

2002 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green